OpenStudy (anonymous):

A jacket usually sells for $60.00. If the jacket is 30% off, and sales tax is 5%, what is the total price of the jacket, including tax?

OpenStudy (anonymous):

$44.10

OpenStudy (king):

55 $

OpenStudy (mathmate):

Work in order of the selling price, and add tax on it. Selling price = regular price *(1-discount) =60*(1-0.30)=$42 Total price = selling price *(1+tax rate) =$42*(1.05) = $44.10